x3rd + x2nd - 2x =2
Answer by Jk22(389)

x^3 + x^2 - 2x -2=0
x^2(x+1)-2(x+1)=0
(x+1)(x^2-2)=0
(x+1)(x-sqrt(2))(x+sqrt(2))=0

=> x=-1, sqrt(2), -sqrt(2)